# Smart Predict–then–Optimize Paradigm for Portfolio Optimization in Real Markets

### v2

@@ 2.1 Portfolio Optimization Modeling
Modern portfolio optimization originates from the mean–variance framework proposed by Markowitz Markowitz [1952], which formulates asset allocation as a trade-off between expected return and risk. Subsequent studies have extended this framework by introducing alternative risk measures, additional constraints, and techniques to address estimation uncertainty, including the Sharpe ratio Sharpe [1998] and coherent risk measures such as Conditional Value-at-Risk (CVaR) Rockafellar and Uryasev [2000]. Robust portfolio optimization further accounts for parameter uncertainty by optimizing portfolio decisions against worst-case scenarios within prescribed uncertainty sets Goldfarb and Iyengar [2003].
From both theoretical and empirical perspectives, Jagannathan and Ma Jagannathan and Ma [2003] demonstrate that portfolio weight constraints can be interpreted as a form of statistical regularization. Such constraints mitigate estimation error and improve out-of-sample risk performance, thereby providing a methodological foundation for incorporating additional constraints and penalty terms in portfolio optimization models.
In particular, transaction fee and portfolio turnover are commonly modeled through penalty terms in the optimization objective to account for realistic trading frictions Boyd et al. [2017].
@@ 2.2 Learning-based Approaches
Reinforcement learning approaches formulate portfolio management as a sequential decision-making problem under a Markov decision process, where an agent learns a policy to maximize cumulative rewards through repeated interaction with a financial environment. A variety of RL Moody and Saffell [2001] and deep reinforcement learning (DRL) Jiang et al. [2017] algorithms have been applied in this context, including DQN Pigorsch and Schäfer [2022], PPO Liu et al. [2024], and DDPG Deng et al. [2017]. While these approaches offer high modeling flexibility, they typically rely on carefully designed reward functions and environment simulators, which may lead to unstable training dynamics and limited interpretability in noisy and non-stationary financial markets Bai et al. [2025].
@@
Within the PtO framework, a wide variety of machine learning and neural network models have been extensively investigated for portfolio optimization. In these approaches, learning models are used to estimate key problem parameters, including expected asset returns, risk measures, or other predictive signals, which are subsequently incorporated into classical portfolio optimization formulations. Representative predictive models include linear regression Deng and Min [2013], support vector machine(SVM) Paiva et al. [2019], tree-based method, and deep neural network (DNN) Freitas et al. [2009], such as recurrent neural network (RNN) and long short-term memory (LSTM) Wang et al. [2020] network. Comprehensive reviews and empirical studies  Krauss et al. [2017], Ma et al. [2021]have examined the effectiveness of these models in portfolio-related applications. By integrating data-driven predictions with optimization-based portfolio construction, PtO methods aim to improve investment performance while maintaining the interpretability and structural advantages of traditional portfolio optimization frameworks.
However, stock market prediction remains a challenging time-series forecasting problem, as financial markets are inherently nonlinear, dynamic, noisy, and chaotic, making direct and accurate prediction difficult Deboeck [1994].
Recent advances in recurrent neural networks such as LSTM Lindemann et al. [2021] and Transformer-based models Wang et al. [2024], Zhang et al. [2023], Nie et al. [2023], Liu et al. [2023] have demonstrated strong effectiveness across a wide range of time-series forecasting tasks. Nevertheless, these architectures often rely on inductive biases such as long-range dependency modeling or periodic pattern extraction, which are particularly suitable for structured or seasonal time series. In contrast, financial time series are characterized by weak periodicity, low signal-to-noise ratios, and frequent regime shifts, limiting the practical gains of highly complex architectures in real-world investment settings.
Moreover, the increased architectural complexity of deep recurrent and Transformer-based models raises concerns regarding training stability and interpretability in high-stakes financial decision-making. For example, recent studies such as DLinear Zeng et al. [2023] question the necessity of complex architectures for time-series forecasting, showing that simple linear models can achieve competitive performance.
@@ 2.3 Smart Predict-then-Optimize (SPO)
The distinction between predict–then–optimize (PtO) and Smart Predict–then–Optimize (SPO) was formally introduced by Elmachtoub and Grigas Elmachtoub and Grigas [2022]. While PtO methods focus on minimizing prediction error independently of the downstream optimization problem, SPO explicitly aligns the learning objective with decision quality. In related literature, this paradigm is also referred to as prediction-focused learning and decision-focused learning Mandi et al. [2024].
@@
This perspective is especially relevant in financial applications. As discussed earlier, stock return prediction is inherently difficult due to market noise, non-stationarity, and complex dynamics, making prediction errors inevitable. At the same time, financial decision-making places stringent requirements on robustness, stability, and risk control, where small prediction errors may lead to disproportionately large economic losses. Under such conditions, learning paradigms that explicitly optimize decision quality rather than pointwise prediction accuracy are particularly valuable. Recent work has further emphasized the importance of robustness in decision-focused learning by proposing regret-based loss functions that explicitly hedge against adverse decision outcomes Schutte et al. [2024], motivating the adoption of robust SPO-based approaches for portfolio optimization.
Preliminary ideas related to this study were explored in an earlier conference version Wang and Hasuike [2025]. The present work substantially extends this line of research by providing a systematic evaluation and more comprehensive modeling. Building upon this initial investigation, the present study conducts a more systematic evaluation of SPO-based methods in portfolio optimization.
In addition, this study adopts PyEPO Tang and Khalil [2024] to implement the SPO loss. PyEPO offers a convenient and flexible framework for SPO modeling and has been increasingly used in recent working papers and preprints.
@@ 3.2 SoftmaxDFL
Motivated by prior work Zhang et al. [2020] that proposes constructing portfolio weights via a softmax layer and directly optimizing financial performance metrics through a loss function, we also consider an alternative learning-based portfolio allocation approach based on a fully differentiable allocation structure. As a differentiable baseline, we consider a neural allocator that directly maps input features to portfolio weights via a softmax layer, denoted as SoftmaxDFL. Unlike SPO-based methods that predict returns and then solve an explicit optimization problem, SoftmaxDFL learns the allocation policy end-to-end without an explicit optimization layer.
@@ 7 Future Work
Several promising directions remain for future research. First, while the proposed framework accounts for decision-aware optimization under predicted returns, incorporating more advanced downside risk control mechanisms—such as distributionally robust optimization and Value-at-Risk (VaR)–type constraints may further enhance robustness against extreme market movements. Second, although the current study adopts simple and stable predictive models, future work may explore the integration of tree-based predictors within the decision-focused learning framework Elmachtoub et al. [2020]. From an empirical finance perspective, tree-based ensemble models have been shown to exhibit strong robustness in noisy return prediction tasks. For example, Krauss et al. Krauss et al. [2017] demonstrate that random forests and gradient-boosted trees achieve competitive and often superior out-of-sample performance compared to deep neural networks in statistical arbitrage settings, where financial signals are characterized by low signal-to-noise ratios. This robustness makes tree-based predictors a promising direction for future integration into decision-focused portfolio optimization frameworks. Finally, despite the inclusion of diversification-promoting mechanisms, achieving an optimal balance between portfolio diversification and return maximization remains an open challenge. Developing principled approaches to jointly model this trade-off within decision-focused portfolio optimization frameworks constitutes an important direction for future research.

### v3

@@ 2.1 Portfolio Optimization Modeling
Modern portfolio optimization originates from the mean–variance framework proposed by Markowitz [18], which formulates asset allocation as a trade-off between expected return and risk. Subsequent studies have extended this framework by introducing alternative risk measures, additional constraints, and techniques to address estimation uncertainty, including the Sharpe ratio [25] and coherent risk measures such as Conditional Value-at-Risk (CVaR) [23]. Robust portfolio optimization further accounts for parameter uncertainty by optimizing portfolio decisions against worst-case scenarios within prescribed uncertainty sets [9].
From both theoretical and empirical perspectives, Jagannathan and Ma [10] demonstrate that portfolio weight constraints can be interpreted as a form of statistical regularization. Such constraints mitigate estimation error and improve out-of-sample risk performance, thereby providing a methodological foundation for incorporating additional constraints and penalty terms in portfolio optimization models.
In particular, transaction fee and portfolio turnover are commonly modeled through penalty terms in the optimization objective to account for realistic trading frictions [2].
@@ 2.2 Learning-based Approaches
Reinforcement learning approaches formulate portfolio management as a sequential decision-making problem under a Markov decision process, where an agent learns a policy to maximize cumulative rewards through repeated interaction with a financial environment. A variety of RL [19] and deep reinforcement learning (DRL) [11] algorithms have been applied in this context, including DQN [22], PPO [14], and DDPG [5]. While these approaches offer high modeling flexibility, they typically rely on carefully designed reward functions and environment simulators, which may lead to unstable training dynamics and limited interpretability in noisy and non-stationary financial markets [1].
@@
Within the PtO framework, a wide variety of machine learning and neural network models have been extensively investigated for portfolio optimization. In these approaches, learning models are used to estimate key problem parameters, including expected asset returns, risk measures, or other predictive signals, which are subsequently incorporated into classical portfolio optimization formulations. Representative predictive models include linear regression [4], support vector machine(SVM) [21], tree-based method, and deep neural network (DNN) [8], such as recurrent neural network (RNN) and long short-term memory (LSTM) [27] network. Comprehensive reviews and empirical studies  [12, 16]have examined the effectiveness of these models in portfolio-related applications. By integrating data-driven predictions with optimization-based portfolio construction, PtO methods aim to improve investment performance while maintaining the interpretability and structural advantages of traditional portfolio optimization frameworks.
However, stock market prediction remains a challenging time-series forecasting problem, as financial markets are inherently nonlinear, dynamic, noisy, and chaotic, making direct and accurate prediction difficult [3].
Recent advances in recurrent neural networks such as LSTM [13] and Transformer-based models [29, 31, 20, 15] have demonstrated strong effectiveness across a wide range of time-series forecasting tasks. Nevertheless, these architectures often rely on inductive biases such as long-range dependency modeling or periodic pattern extraction, which are particularly suitable for structured or seasonal time series. In contrast, financial time series are characterized by weak periodicity, low signal-to-noise ratios, and frequent regime shifts, limiting the practical gains of highly complex architectures in real-world investment settings.
Moreover, the increased architectural complexity of deep recurrent and Transformer-based models raises concerns regarding training stability and interpretability in high-stakes financial decision-making. For example, recent studies such as DLinear [30] question the necessity of complex architectures for time-series forecasting, showing that simple linear models can achieve competitive performance.
@@ 2.3 Smart Predict-then-Optimize (SPO)
The distinction between predict–then–optimize (PtO) and Smart Predict–then–Optimize (SPO) was formally introduced by Elmachtoub and Grigas [6]. While PtO methods focus on minimizing prediction error independently of the downstream optimization problem, SPO explicitly aligns the learning objective with decision quality. In related literature, this paradigm is also referred to as prediction-focused learning and decision-focused learning [17].
@@
This perspective is especially relevant in financial applications. As discussed earlier, stock return prediction is inherently difficult due to market noise, non-stationarity, and complex dynamics, making prediction errors inevitable. At the same time, financial decision-making places stringent requirements on robustness, stability, and risk control, where small prediction errors may lead to disproportionately large economic losses. Under such conditions, learning paradigms that explicitly optimize decision quality rather than pointwise prediction accuracy are particularly valuable. Recent work has further emphasized the importance of robustness in decision-focused learning by proposing regret-based loss functions that explicitly hedge against adverse decision outcomes [24], motivating the adoption of robust SPO-based approaches for portfolio optimization.
Preliminary ideas related to this study were explored in an earlier conference version [28]. The present work substantially extends this line of research by providing a systematic evaluation and more comprehensive modeling. Building upon this initial investigation, the present study conducts a more systematic evaluation of SPO-based methods in portfolio optimization.
In addition, this study adopts PyEPO [26] to implement the SPO loss. PyEPO offers a convenient and flexible framework for SPO modeling and has been increasingly used in recent working papers and preprints.
@@ 3.2 SoftmaxDFL
Motivated by prior work [32] that proposes constructing portfolio weights via a softmax layer and directly optimizing financial performance metrics through a loss function, we also consider an alternative learning-based portfolio allocation approach based on a fully differentiable allocation structure. As a differentiable baseline, we consider a neural allocator that directly maps input features to portfolio weights via a softmax layer, denoted as SoftmaxDFL. Unlike SPO-based methods that predict returns and then solve an explicit optimization problem, SoftmaxDFL learns the allocation policy end-to-end without an explicit optimization layer.
@@ 7 Future Work
Several promising directions remain for future research. First, while the proposed framework accounts for decision-aware optimization under predicted returns, incorporating more advanced downside risk control mechanisms—such as distributionally robust optimization and Value-at-Risk (VaR)–type constraints may further enhance robustness against extreme market movements. Second, although the current study adopts simple and stable predictive models, future work may explore the integration of tree-based predictors within the decision-focused learning framework [7]. From an empirical finance perspective, tree-based ensemble models have been shown to exhibit strong robustness in noisy return prediction tasks. For example, Krauss et al. [12] demonstrate that random forests and gradient-boosted trees achieve competitive and often superior out-of-sample performance compared to deep neural networks in statistical arbitrage settings, where financial signals are characterized by low signal-to-noise ratios. This robustness makes tree-based predictors a promising direction for future integration into decision-focused portfolio optimization frameworks. Finally, despite the inclusion of diversification-promoting mechanisms, achieving an optimal balance between portfolio diversification and return maximization remains an open challenge. Developing principled approaches to jointly model this trade-off within decision-focused portfolio optimization frameworks constitutes an important direction for future research.